FAX Cover Page

FAX Page Count: 2

17 March 2000, 9:42 A.M.

To:

Pyotr Ilyitch Tchyfontski Research Associate Professor Division of Fontal Conversion St. Leninberg University of the New State Telephone: Telephone: Telephone: Telephone: FAX:

From:

Ludwig von Beetfonten # 5 Fifth Ave. New Font City, New Fontiac United States of Fontagraphy 12345-6789 Telephone: 030-1234567 Telephone: Telephone: Telephone: FAX: 030-55535552

The Font Institute

We Do Fonts Right

March 17, 2000

# 5 Fifth Ave. New Font City, New Fontiac United States of Fontagraphy 12345-6789

Pyotr Ilyitch Tchyfontski Research Associate Professor Division of Fontal Conversion St. Leninberg University of the New State

Tovarich Pyotr:

In terms of the relationships between font descriptions and musical typography, it seems to me that only in the case of the early 1800s should we think of such a really odd notion.

Yours in fontal health,

Ludwig von Beetfonten Research Associate Professor Division of Font Music Insertions Vienna University